Consider ABC an equilateral triangle of sides 4cm.

(a) determine the position of the center of gravity of the equilateral triangle

(b) determine and construct the set of points C such that IMA + MB + MCII = 12/√3

(c) name the set of points C



1
Expert's answer
2021-09-28T15:18:16-0400

a.



The position of the center of mass is "(\frac{l}{2},x)." Wher l is the length of the sides of the triangle and x is gotten as follows.

"\frac{2}{2\sqrt3}=\frac{x}{2}\\
x=\frac{2}{3}\sqrt3"

Therefore, the position of the center of mass is "(2,\frac{2}{3}\sqrt3)" .

b.

"|MA|=2\sqrt3-\frac{2}{3}\sqrt3=\frac{4}{3}\sqrt3\\
|MB|=2\sqrt3-\frac{2}{3}\sqrt3=\frac{4}{3}\sqrt3\\
|MA+MB+MC|=\frac{12}{\sqrt3}\\
\frac{4}{3}\sqrt3+\frac{4}{3}\sqrt3+|MC|=4\sqrt3\\
|MC|=\frac{4}{3}\sqrt3"

Therefore, the set of points of C is "(4,0)" . Since

"MA=\sqrt{(2-4)^2+(\frac{2}{3}\sqrt3-0)^2}=\frac{4}{3}\sqrt3"



c.

the set of point of C is "(4,0)"